# KeySpace: Public Key Infrastructure Considerations in Interplanetary Networks

### v1

@@ Abstract
As satellite networks grow larger and begin to incorporate interplanetary communication, there is an increasing interest in the unsolved problem of how to approach PKI in these conditions. In this paper we explore the goals and requirements for implementing key management systems in satellite networks, focusing on megaconstellations and interplanetary networks. We design a set of standardized experiments which can be used to compare systems against one another for particular network topologies. Using these, we demonstrate that terrestrial PKI techniques are feasible in highly distributed interplanetary networks, showing that it is possible to configure PKI systems to achieve efficient low-latency connection establishment, and minimize the impact of attacks through effective revocations.
We evaluate this by building the Deep Space Network Simulator (DSNS), a novel network simulator aimed at efficient simulation of large space networks. We run simulations evaluating connection establishment and key revocation under a wide range of PKI configurations. Finally, we propose and evaluate two additional configuration options: OCSP Hybrid, and the use of relay nodes as a firewall. Together these minimize the extent of the network an attacker can reach with a compromised key, and reduce the attacker’s load on interplanetary relay links.
I Motivation
@@
This results in a network topology that has a number of sporadic, high-latency links (see Figure 1). The network naturally partitions itself based on these long-distance links, with each partition being (for the most part) connected. This type of network is therefore considered to be a Delay-Tolerant Network (DTN), an architecture which has been explored extensively in related works. In DTNs, key management is considered to be an unsolved problem [templinDelay2014] and is an active area of research, with the majority of research focusing on the unpredictable nature of the network, making use of techniques like gossip protocols and web of trust to provide probabilistic assurances of success despite the network’s unpredictable movement and connectivity [khareWeaving1997, garfinkelPGP1995, eschenauerKeyManagement2002, menesidouCryptographic2017, deandradeFully2016].
@@
Contributions
In this paper, we outline the goals and requirements for deploying and operating PKI in large satellite networks, with a particular focus on interplanetary networks with long-distance, intermittent relays. We construct a standardized set of experiments for comparing different key management systems to one another for a given network topology, looking in particular at the performance of connection establishment procedures, and the speed at which key revocations and updates can occur so that the damage caused by a compromised key can be minimized. We evaluate these through the use of simulations, building the Deep Space Network Simulator (DSNS): an extensible system for the realistic simulation of large networks of satellites. Through this analysis we demonstrate, for the first time, that commonly used terrestrial PKI protocols can be employed in satellite networks. We also establish configurations which optimize performance and security, and propose novel variations on these protocols which provide more effective revocations to protect remote segments more quickly.
II Background
@@
DTNs encompass a wide range of networks including mobile ad-hoc networks, wireless sensor networks, and networks in space [torgersonDelayTolerant2007]. These networks are characterized by their lack of persistent connectivity and unpredictable network topology. This requires changes to the network stack, particularly in routing: end-to-end routes cannot always be computed ahead of time, or even established in the first place, due to the structure of the network. Instead, standards like the Bundle Protocol (BP) are used [scottBundle2007]. This protocol encapsulates data in “bundles”, with each bundle containing data for one or more applications. These bundles are sent around the network in a store-and-forward manner, and can be sent over multiple paths to maximize the chances of reaching their destination.
Much like nodes in a DTN, satellites communicate with one another using radio or optical links. However, the crucial difference here is that connections between satellites can be predicted ahead of time nearly perfectly, since their movement is fixed. As a result, some of the design considerations involved in BP can be disregarded – instead of sending messages over multiple paths to maximize the chance it reaches its destination, we can compute the optimal route ahead of time. However, these networks can still make use of the store-and-forward aspects of BP’s message delivery. Sometimes referred to as the Interplanetary Internet (IPN), there also exists a body of work addressing routing and other aspects of networking under these particular constraints [akyildizState2004, elalaouiMARS2020, alhilalSky2019]. The Internet Society’s Interplanetary Chapter has also produced a report with recommendations for the architecture and governance of such a network. [ipnsigSolar2023].
Public-Key Infrastructure (PKI)
@@
At a high level, a PKI attaches identities to public-private keypairs to secure communication in larger networks. This is usually achieved with certificates provided by Certificate Authorities (CAs), with a cryptographic signature asserting the legitimacy of a keypair. A network participant can thus check that a keypair is legitimate – all they need is the certificate for the keypair, signed by a trusted CA. A root CA typically delegates the ability to issue further certificates to other CAs, resulting in a signature chain. This system is near-universal in the terrestrial internet, using the X.509 standard [boeyenInternet2008].
If a key has been compromised or otherwise needs to exit circulation, a revocation is issued, stating that the certificate is no longer valid. Therefore, in order to trust a certificate the recipient must also verify that it has not been revoked. One way to achieve this is through Certificate Revocation Lists (CRLs): the CA periodically produces a list of revoked certificates, which can be broadcast or requested by users in the network. Alternatively, the Online Certificate Status Protocol (OCSP) may be used [santesson5092013]. This protocol allows recipients to query CAs about the status of certificates when receiving a message, removing the requirement for large CRLs to be distributed. A variant known as “OCSP Stapling” requires senders to instead query the CA about their own certificates status and attach the response to the message, thus reducing the burden on recipients [eastlake3rdTransport2011]. This places the burden of validating a certificate on the node initiating communication, reducing the risk of denial-of-service attacks in which an attacker makes a large number of message requests. It also allows the message exchange with a CA to be performed ahead of time, allowing communication over long-distance links to occur as soon as a link becomes available.
@@
However, in DTNs both CRLs and OCSP are largely dismissed as infeasible: CRLs are too bandwidth-intensive and OCSP requires constant low-latency access to CAs, which cannot always be guaranteed. Instead, DTNs often make use of gossip protocols for key distribution – in a decentralized network with unpredictable connectivity, nodes exchange information with their peers on key validity and revocations to ensure this information spreads as far as possible [eschenauerKeyManagement2002]. In this paper we demonstrate that these assumptions do not hold for all DTNs, and that the predictable links and network topology of space networks enables many terrestrial PKI concepts to be feasible. Furthermore, once the above arguments have been addressed it can be argued that PKI is ideal for space applications. The protocols scale particularly well to huge numbers of nodes, thus future proofing space networks as they continue to grow even larger.
Current Systems
@@
Many satellite systems do not use PKI, opting instead to use pre-shared encryption keys. This is significantly simpler to deploy in systems consisting of only a single satellite, as the key can be loaded onto the satellite before launch, but does not provide perfect forward secrecy. It also presents additional risks in case keys are leaked (see the GK-2A weather satellite [xrit-rx]), or if the key is broken or reverse engineered (see the COMS-1 meteorological satellite [lrit-key-dec]). Some systems support over-the-air rekeying in this case, using more tightly controlled keys to authenticate the operation of replacing session keys [andersonAuthentication2023, ccsdsSpace2020].
@@
The majority of current satellite constellations are privately owned (e.g., Starlink), so less is known about their internal operation. The Iridium satellite consteallation uses the GSM encryption algorithm with a pre-shared key stored in the SIM card [veenemanIridium2021], and has no mechanism for updating keys short of rewriting or replacing the SIM. The key management of the satellites themselves is not known.
III Related Work
@@
Internet Drafts
The most recent proposed standards for the Bundle Protocol include the BPSec security extensions [burleighBundle2022, birraneBundle2022], which add options for integrity and confidentiality to secure BP communication. Notably, the BPSec standard does not recommend a specific method for key management; instead, “BPSec assumes that key management is handled as a separate part of network management” [birraneBundle2022]. In contrast, in this paper we focus on the key management mechanism, assuming the underlying cryptosystem to be provided through the use of BPSec or a similar protocol.
@@
In [ccsdsSpace2011], the Consultative Committee for Space Data Systems (CCSDS) give a high-level overview of many potential key management concepts in space DTNs. There is also an ongoing CCSDS project to design a cloud-based CA for satellite systems, to be used as a common route of trust particularly in shared systems [ccsdsProject2021]. Results from our simulations will help to inform this and other projects – this is discussed further in Section VIII.
@@
Some works such as [johnsonProviding2013] adapt PKI for use in satellite networks – their system is very similar to OCSP stapling systems discussed above, with nodes requesting an authentication pass from a CA ahead of communication. On the other hand, [sethPractical2005] argues that PKI is not suitable for use in DTNs, instead suggesting the use of Hierarchical Identity Based Cryptography (HIBC), in which keys are derived from public information about nodes, and a hierarchy of authorities manages keys. This is argued to be better than traditional PKI, as the public parameters do not need to be verified as frequently [asokanSecuring2007], at the cost of requiring additional secrets to be distributed out of channel. We argue that this is not necessary for satellite networks, as each network segment is sufficiently connected that it is always possible to query a trusted authority, provided there are sufficiently many distributed throughout the network. We demonstrate this using simulations in Section VII. HIBC also raises concerns about large-scale deployment; distributing secrets and maintaining identities out-of-channel becomes significantly more difficult as the number of users grows.
@@
IV Threat Model
In this work we consider two main scenarios. The first of these is benign communication: nodes in the network wish to communicate with one another, which requires the validity of the sender’s certificate to be checked. This adds a time delay before communication can occur, which can be quite significant in satellite networks with long-distance links. This is a particular concern when communication occurs over a sporadic link between network segments; any time spent establishing the sender’s identity is time in which commmunication cannot occur, taking away from the total duration of the link’s uptime. We aim to minimize this time overhead, while also optimizing for latency and message loss due to timeout.
In the second scenario, an attacker compromises the key of a legitimate user on the network, aiming to use it for malicious operation or communication. There are a number of ways this could occur: compromising a satellite (e.g., by compromising telecommand authentication, taking control of a satellite via a compromised ground segment, or via a malicious hosted payload), intercepting improperly encrypted messages, compromising the supply chain, etc. We can expect that in a federated satellite network with many operators and users this is likely to happen occasionally due to the number of keys involved, so it is critical that revocations are handled as effectively as possible. The long distances between nodes in an interplanetary network and inherent speed-of-light limit means there may be no way to entirely prevent the attacker from doing damage; instead, our aim is to minimize the extent of the damage from the moment the attack is identified, and a revocation issued.
The segmentation of the network further partitions this attack scenario: we must consider what happens if the attacker is present in the network segment from which the revocation originates, or in a different network segment. In the latter case the attacker is likely to be able to communicate within their own segment of the network for a reasonable amount of time, since the revocation messages will take a long time to traverse the interplanetary links. Our main concern in this case is securing the attacker’s segment as early as possible, and ensuring their messages cannot spread to further segments.
V Framework Design
In this section we establish the goals and requirements for key management in satellite networks, constructing from them a set of standardized experiments which can be used to compare protocols against one another. Rather than being specific to one network topology or configuration, these are instead designed to be high-level tests that can be used on any network to compare different protocols side by side. This allows network operators to test protocols in the context of the specifics of their network, enabling well-informed decision making about which protocol and configuration to choose, balancing performance against security. In Section VI we design experiments using standard terrestrial PKI protocols on interplanetary satellite networks, proposing and testing improvements specific to the unique topology of these networks.
@@ V Framework Design
  • Intercompatible: We argue that in order for interplanetary connectivity to be effective, protocols will need to work alongside existing internet infrastructure. Protocols may be designed to work seamlessly with existing internet PKI, or utilize a translation layer. Protocols should also be able to be made compatible with BPSec [birraneBundle2022], which provides mechanisms to secure communication but leaves key management as an open question.
V-A Scenarios
@@
V-A1 Connection Establishment
In the first scenario two nodes want to communicate with one another, but must first establish a connection and verify the other’s identity.
In order to eliminate the need for traffic modeling we consider the connection establishment process between every pair of communicating nodes in the network, giving a distribution of values for each metric. We focus on the following metrics:
  • Dropped messages (%)
@@ V-A2 Key Revocation
Coverage time $T_{R}(s)$ is assessed for each major network segment by finding the time at which the relevant CA has received the revocation message, adjusting to account for query time, and finding the time at which all nodes are covered. As a result, for certain segments the coverage time may be negative – if a revocation reaches the CA quickly, but it takes a long time for the attacker’s message to reach that same CA (for instance, if there is a long-distance link between them), then nodes under that CA may be protected even from messages initialized before the revocation was started. Also note that in some cases such as OCSP Stapling the coverage will depend on the attacker’s location and CA; in these cases, results are gathered for every possible attacker location and aggregated by segment. We also assess attack penetration $P(s)$ for cases where certificate validity information is cached, with a fixed attacker location chosen to be in an opposite orbital shell from the CA in segment $s$.
VI Experiment Design
@@
VI-A Key Management Systems
In our experiments we evaluate a number of PKI systems based on the existing and proposed mechanisms discussed in Section III. In each case, nodes have a public/private keypair which is used to sign and encrypt messages. These keypairs are accompanied by a certificate – that is, a chain of signatures tracing back to a Certificate Authority (CA) vouching for its legitimacy. We specifically consider the following systems to verify a certificate’s validity, ensuring it has not been revoked:
  • Certificate Revocation Lists (CRLs): CAs each maintain a list of revoked certificates, and when a node needs to check the validity of a certificate it requests the CRL from its CA. Any certificates on this list are rejected, and others are accepted provided there is a trusted CA in the certificate chain. CRLs can be cached for a configurable amount of time – after this time has elapsed, the node will request a new CRL on the next message received. To reduce size, CRLs can be built using hash tables [bhuttaPublic2017] or using delta CRLs [cooperMore2000].
  • CRL Broadcast: Similar to the above, but CAs broadcast CRLs at regular intervals, instead of providing them on request.
  • Online Certificate Status Protocol (OCSP): On receipt of a signed message, nodes send a message to their CA with a request for the certificate’s validity information. The response can be cached to reduce the latency of the validation of subsequent messages.
  • OCSP Stapling: Similar to the above, CAs provide up-to-date validity information on certificates. However, in this case the node sending the message requests its CA for a signed message with validity information. This message is sent alongside (i.e., stapled to) the original message, reducing the burden on the recipient to validate certificates.
  • OCSP with Validators: Proposed in [viswanathanArchitecture2016], this method is similar to OCSP stapling, but messages are instead sent via a CA, which can staple the certificate validity to the message in-transit. This removes the need for a full handshake in certificate validation, and can significantly reduce latency if the CA is en route to the message destination. This shares some similarities with the proposed “Revocation in the Middle” technique, using middleboxes to provide revocation information [szalachowskiRITM2016].

### v3

@@ Abstract
As satellite networks expand to encompass megaconstellations and interplanetary communication, the need for effective Public Key Infrastructure (PKI) becomes increasingly pressing. This paper addresses the challenge of implementing PKI in these complex networks, identifying the essential goals and requirements.
We develop a standardized framework for comparing PKI systems across various network topologies, enabling the evaluation of their performance and security. Our results demonstrate that terrestrial PKI techniques can be adapted for use in highly distributed interplanetary networks, achieving efficient low-latency connection establishment and minimizing the impact of attacks through effective revocation mechanisms. This result has significant implications for the design of future satellite networks, as it enables the reuse of existing PKI solutions to provide increased compatibility with terrestrial networks.
We evaluate this by building the Deep Space Network Simulator (DSNS), a novel tool for efficiently simulating large space networks. Using DSNS, we conduct comprehensive simulations of connection establishment and key revocation under a range of network topologies and PKI configurations. Furthermore, we propose and evaluate two new configuration options: OCSP Hybrid, and the use of relay nodes as a firewall. Together these minimize the extent of the network an attacker can reach with a compromised key, and reduce the attacker’s load on interplanetary relay links.
1 Motivation
@@
This results in a network topology that has a number of sporadic, high-latency links (see Figure 1). The network naturally partitions itself based on these long-distance links, with good connections within each partition. This type of network is considered to be a Delay-Tolerant Network (DTN), an architecture which has been explored extensively in related works. In DTNs, key management is considered to be an unsolved problem [templinDelay2014] and is an active area of research, with the majority of research focusing on the unpredictable nature of the network, making use of techniques like gossip protocols and web of trust to provide probabilistic assurances of success despite the network’s unpredictable movement and connectivity [khareWeaving1997, garfinkelPGP1995, eschenauerKeyManagement2002, menesidouCryptographic2017, deandradeFully2016].
@@
Contributions. In this paper, we outline the goals and requirements for deploying and operating PKI in large satellite networks, with a particular focus on interplanetary networks with long-distance, intermittent relays. We construct a standardized set of experiments for comparing different key management systems to one another for a given network topology, looking in particular at the performance of connection establishment procedures, and the speed at which key revocations and updates can occur so that the damage caused by a compromised key can be minimized. We evaluate these through the use of simulations, building the Deep Space Network Simulator (DSNS): an extensible system for the realistic simulation of large networks of satellites. Through this analysis we demonstrate, for the first time, that commonly used terrestrial PKI protocols can be employed in satellite networks. We also establish configurations which optimize performance and security, and propose novel variations on these protocols which provide more effective revocations to protect remote segments more quickly. We do not compare to novel protocols designed for DTNs; rather, we demonstrate that existing protocols can be effectively used in satellite networks, thereby maintaining compatibility with terrestrial networks.
2 Background
@@
Delay Tolerant Networks (DTNs). DTNs encompass a wide range of networks including mobile ad-hoc networks, wireless sensor networks, and networks in space [torgersonDelayTolerant2007]. These networks are characterized by a lack of persistent connectivity and unpredictable network topology. This requires changes to the network stack, particularly in routing: end-to-end routes cannot always be precomputed, or even established in the first place, due to the structure of the network. Instead, standards like the Bundle Protocol (BP) are used, enabling store-and-forward delivery [scottBundle2007], and proposed routing techniques often rely upon probabilistic methods [grasicEvolution2011, leguayDTN2005].
However, unlike other types of DTN the movement of satellites is fixed, so connectivity can be predicted ahead of time with near perfect precision. As a result, optimal routes can be computed ahead of time using techniques like Contact Graph Routing (CGR) [aranitiContact2015]; however, store-and-forward delivery is still required due to the sporadic nature of links. Sometimes referred to as the Interplanetary Internet (IPN), there exists a good body of work addressing networking under these particular constraints [akyildizState2004, elalaouiMARS2020, alhilalSky2019], including recommendations for architecture and governance from the Internet Society [ipnsigSolar2023], and protocol and architecture recommendations from the Consultative Committee for Space Data Systems (CCSDS) [ccsdsReference2008, ccsdsOverview2014].
@@
Public-Key Infrastructure (PKI). A PKI attaches identities to public-private keypairs to secure communication in larger networks. It does this through certificates provided by Certificate Authorities (CAs), which verify a keypair’s identity. A network participant can check a keypair’s legitimacy by verifying its certificate, signed by a trusted CA which is part of a signature chain established by the root CA. This system is near-universal in the terrestrial internet, using the X.509 standard [boeyenInternet2008].
To verify a certificate’s trustworthiness, the recipient must check that it has not been revoked. This can be achieved using Certificate Revocation Lists (CRLs), listing all revoked certificates, or using the Online Certificate Status Protocol (OCSP), which allows devices to query the CA about the status of a certificate [santesson5092013]. A variant called OCSP Stapling shifts the burden to the sender, who queries the CA and attaches the response to the message, reducing the risk of denial-of-service attacks and allowing for faster communication over long-distance links [eastlake3rdTransport2011].
@@
However, in DTNs both CRLs and OCSP are largely dismissed as infeasible: CRLs are too bandwidth-intensive and OCSP requires constant low-latency access to CAs, which cannot always be guaranteed. Instead, DTNs often make use of gossip protocols for key distribution – nodes exchange information with their peers in the hope that this information spreads as far as possible [eschenauerKeyManagement2002]. In this paper we demonstrate that these assumptions do not hold for all DTNs, and that the predictable links and network topology of interplanetary networks enables many terrestrial PKI concepts to be effective. Furthermore, once the above arguments have been addressed it can be argued that PKI is ideal for space applications, scaling particularly well to huge numbers of nodes, thus future-proofing space networks as they continue to grow even larger.
@@
Current Systems. Many current satellite systems opt to use pre-shared encryption keys. This is significantly simpler to deploy in small systems, as keys can be loaded onto the satellite before launch, but does not scale well to systems with more satellites or shared operation. It also presents additional risks in cases where keys are leaked, broken, or reverse engineered (see the GK-2A and COMS-1 satellites [xrit-rx, lrit-key-dec]). Some systems support over-the-air rekeying in this case, using tightly controlled keys to authenticate the operation of replacing session keys [andersonAuthentication2023, ccsdsSpace2020].
@@
The majority of current satellite constellations are privately owned (e.g., Starlink), so less is known about their internal operation. The Iridium constellation uses pre-shared keys stored in the SIM card, with no mechanism for updating keys short of rewriting or replacing the SIM [veenemanIridium2021]. The key management of the satellites themselves is not known.
3 Related Work
@@
3.1 Internet Drafts
The most recent proposed standards for the Bundle Protocol include the BPSec security extensions [burleighBundle2022, birraneBundle2022], which add options for integrity and confidentiality to secure BP communication. Notably, this standard does not recommend a specific method for key management, assuming it to be handled separately [birraneBundle2022]. In contrast, in this paper we focus on the key management mechanism, assuming the underlying cryptosystem to be provided through BPSec or a similar protocol.
@@ 3.1 Internet Drafts
In [ccsdsSpace2011], the Consultative Committee for Space Data Systems (CCSDS) give a high-level overview of many potential key management concepts in space DTNs. There is also an ongoing CCSDS project to design a cloud-based CA for satellite systems, to be used as a common route of trust particularly in shared systems [ccsdsProject2021]. Results from our simulations will help to inform this and other projects – this is discussed further in Section 8.
@@ 3.2 Academic Works
Some works such as [johnsonProviding2013] adapt PKI for use in satellite networks – their system is very similar to OCSP stapling systems discussed above, with nodes requesting an authentication pass from a CA ahead of communication. On the other hand, [sethPractical2005] argues that PKI is not suitable for use in DTNs, instead suggesting the use of Hierarchical Identity Based Cryptography (HIBC), in which keys are derived from public information about nodes, and a hierarchy of authorities manages keys. This is argued to be better than traditional PKI, as the public parameters do not need to be verified as frequently [asokanSecuring2007], at the cost of requiring additional secrets to be distributed out of channel. We argue that this is not necessary for satellite networks, as each network segment is sufficiently connected that it is always possible to query a trusted authority, provided there are sufficiently many distributed throughout the network. We demonstrate this using simulations in Section 7. HIBC also raises concerns about large-scale deployment; distributing secrets and maintaining identities out-of-channel becomes significantly more difficult as the number of users grows.
@@ 3.3 Simulators
4 Threat Model
In this paper we consider an attacker that has gained access to the private key of a legitimate user on the network, and has at least one node in the network under their control. Using this key they can craft messages as if they were the legitimate user. There are a number of ways this could occur: compromising a satellite (e.g., by compromising telecommand authentication, taking control of a satellite via a compromised ground segment, or via a malicious hosted payload), intercepting improperly encrypted messages, compromising the supply chain, etc.
At a high level, the goal of the attacker is to cause damage within the network by using the compromised key to communicate with as many nodes as possible, for as long as possible. This will look different depending on the nature of the network and the compromised device: for example, producing forged measurement data, impersonating users in communication, or forging telecommand packets to damage the device itself.
When the attacker is discovered, the compromised key must be revoked. The long distances between nodes in an interplanetary network and inherent speed-of-light limit means the goal is not to prevent the attacker entirely, but instead to minimize the extent of the damage from the moment the attack is identified. This is achieved by propagating the revocation as efficiently as possible.
5 Framework Design
In this section we establish the goals and requirements for key management in satellite networks, constructing from them a set of standardized experiments which can be used to compare protocols against one another. Rather than being limited to one network topology or configuration, these are instead designed to be high-level tests that can be used on any network to compare different protocols side by side. This allows network operators to test protocols in the context of their own network configuration, enabling well-informed decision making about which configuration to choose, balancing performance against security. In Section 6 we design experiments using standard terrestrial PKI protocols in interplanetary satellite networks, and evaluate improvements tailored to the unique topology of these networks.
@@ 5 Framework Design
  • Intercompatible: We argue that in order for interplanetary connectivity to be effective, protocols will need to work alongside existing internet infrastructure. Protocols may be designed to work seamlessly with existing protocols, or utilize a translation layer. Protocols should also be able to be made compatible with BPSec [birraneBundle2022], which provides mechanisms to secure communication but leaves key management as an open question.
Our focus is on adapting PKI mechanisms to the environment of interplanetary networks, rather than on the underlying cryptosystem itself. While considerations such as storage, computational requirements, and bandwidth usage are crucial in this context, they are inherently tied to the choice of cryptosystem and are therefore outside the scope of this work. By concentrating on the feasibility of applying terrestrial PKI mechanisms in space, we establish a foundation for future research, which can then focus on optimizing the cryptosystem and its components to meet the unique demands of interplanetary networks.
5.1 Scenarios
@@
5.1.1 Connection Establishment
In the first scenario one node wants to communicate with another, but must first prove its identity. To do so it sends a signed message with a certificate attached. The recipient validates the certificate via validity information either stapled to the message or requested from a CA.
This adds a time delay before communication can occur, which can be quite significant in satellite networks with long-distance links. This is a particular concern when communication occurs over a sporadic link between network segments; any time spent establishing the sender’s identity is time in which commmunication cannot occur, taking away from the total duration of the link’s uptime.
@@ 5.1.2 Key Revocation
Coverage time $T_{R}(s)$ (for segment $s$) is assessed relative to the revocation time $t_{0}$ by finding the time at which each segment’s CA has received the revocation message, adjusting to account for query time, and taking the greatest such time within the segment. As a result, for certain segments the coverage time may be negative – if a revocation reaches the CA quickly, but it takes a long time for the attacker’s message to reach that same CA (for instance, if there is a long-distance link between them), then nodes under that CA may be protected even from messages initialized before the revocation was started. Also note that in some cases such as OCSP Stapling the coverage will depend on the attacker’s location/CA; in these cases, results are gathered for every possible attacker location and aggregated by segment. We also assess attack penetration $P(s)$ for cases where certificate validity information is cached, with a fixed attacker location chosen to be in an opposite orbital shell from the CA in segment $s$.
6 Experiment Design
@@
6.1 Key Management Systems
In our experiments we evaluate a number of PKI systems based on the existing and proposed mechanisms discussed in Section 3. In each case, nodes have a public/private keypair which is used to sign and encrypt messages. These keypairs are accompanied by a certificate – that is, a chain of signatures tracing back to a Certificate Authority (CA) vouching for its legitimacy. We specifically consider the following systems to verify a certificate’s validity, ensuring it has not been revoked:
  • Certificate Revocation Lists (CRLs): CAs each maintain a list of revoked certificates, which are sent to nodes on request. Any certificates on this list are rejected, and others are accepted provided there is a trusted CA in the certificate chain. CRLs are cached, with a new CRL requested on the next message received after expiry. To reduce size, CRLs can be built using hash tables [bhuttaPublic2017] or using delta CRLs [cooperMore2000].
  • CRL Broadcast: Similar to the above, but CAs broadcast CRLs at regular intervals, instead of on request.
  • Online Certificate Status Protocol (OCSP): On receipt of a signed message, nodes request the certificate’s validity information from their CA. The response is cached to reduce the overhead on subsequent messages.
  • OCSP Stapling: Similar to the above, but the node sending the message requests its CA for a signed message with validity information. This is sent alongside the original message, reducing the burden on the recipient.
  • OCSP with Validators: Proposed in [viswanathanArchitecture2016], messages are instead sent via a CA, which staples the certificate validity to the message in-transit. This removes the need for a full handshake in certificate validation, and can significantly reduce latency if the CA is en route to the message destination. This shares some similarities with the proposed “Revocation in the Middle” technique, using middleboxes to provide revocation information [szalachowskiRITM2016].